# Initialization Matters for Adversarial Transfer Learning

### v1

@@ Abstract
With the prevalence of the Pretraining-Finetuning paradigm in transfer learning, the robustness of downstream tasks has become a critical concern. In this work, we delve into adversarial robustness in transfer learning and reveal the critical role of initialization, including both the pretrained model and the linear head. First, we discover the necessity of an adversarially robust pretrained model. Specifically, we reveal that with a standard pretrained model, Parameter-Efficient Finetuning (PEFT) methods either fail to be adversarially robust or continue to exhibit significantly degraded adversarial robustness on downstream tasks, even with adversarial training during finetuning. Leveraging a robust pretrained model, surprisingly, we observe that a simple linear probing can outperform full finetuning and other PEFT methods with random initialization on certain datasets. We further identify that linear probing excels in preserving robustness from the robust pretraining. Based on this, we propose Robust Linear Initialization (RoLI) for adversarial finetuning, which initializes the linear head with the weights obtained by adversarial linear probing to maximally inherit the robustness from pretraining. Across five different image classification datasets, we demonstrate the effectiveness of RoLI and achieve new state-of-the-art results.
@@ 1 Introduction
In this work, we comprehensively study six popular finetuning methods in adversarial transfer learning. We discover that initialization, including 1) the backbone initialization, i.e., pretrained model, and 2) the linear head which adapts features to the target domain, plays a critical role in adversarial transfer learning. We start by comparing the performance of adversarial finetuning when initialized with either a robust or a standard pretrained model. Surprisingly, all PEFT methods fail or exhibit significantly inferior performance when initialized with a standard pretrained model compared to being initialized with a robust pretrained model. Furthermore, we observe that a large-scale pretrained model, such as CLIP [37], cannot alleviate this phenomenon, highlighting the critical role of robust pretraining in adversarial transfer learning.
Given a robust pretrained model, surprisingly, we discover that adversarial linear probing outperforms other methods on certain datasets, e.g., Caltech256 and Stanford Dogs. We further demonstrate that this is mainly because linear probing excels in inheriting robustness from pretraining compared to other methods. Moreover, we identify the ability to adapt features to the specific target domain as a key factor influencing linear probing’s success, supported by a strong correlation between transferred accuracy and robustness. Based on these insights, we introduce Robust Linear Initialization (RoLI), which initializes the model’s linear head with the weights obtained from adversarial linear probing. With RoLI followed by adversarial finetuning methods, we maximize the inherited robustness along with a strong feature adaptation ability, resulting in improved performance. In summary, our contributions are as follows:
@@ 4 Robust Pretrained Model Matters
Figure 4: A robust pretrained model exhibits more semantically structured gradients compared to a standard pretrained model. Visualizations are obtained by computing the loss gradient w.r.t the input image on Caltech256.
@@ 5 Initialization Matters For Finetuning
Table 1: PEFT methods, excluding Linear, demonstrate strong performance on average across CIFAR10 and Caltech256 datasets. With adversarial finetuning, PEFT methods outperform Full-FT and Linear with random linear initialization.
@@ 5.1 Random Linear Initialization
To answer this question, we first evaluate the adversarial robustness of six adversarial finetuning techniques. To adapt the robust pretrained model to the downstream tasks, we initialize the linear head with Random Linear Initialization (RanLI) following [42]. We perform adversarial finetuning across two model architectures, e.g., ViT and Swin, on the CIFAR10 and Caltech256 datasets. Additional details regarding the training process can be found in Sec. 5.4 and in the supplementary materials.
@@ 5.4 Robust Linear Initialization (RoLI)
In our experiments, we use five datasets: CIFAR10 [26], CIFAR100 [26], Caltech256 [15], Caltech-UCSD Birds-200-2011 (CUB200) [45], and Stanford Dogs (Dogs) [23]. For low-resolution image datasets like CIFAR10 and CIFAR100, we resize the image to match the model input shape. For high-resolution image datasets, we apply random resizing and cropping in training and center crop in testing. It’s important to note that we integrate the resizing and normalization process directly into the model since the attacker does not have access to pre-processed images. Additional details are provided in the supplementary material.
@@ Methods.
Training Details.
We use the same optimizer, AdamW [33], and a cosine scheduler with warm-up for all finetuning methods. To determine the optimal learning rate and weight decay values, we conduct a grid search, and the ranges of hyper-parameters along with the optimal combinations can be found in the supplementary material. We keep method-specific hyper-parameters constant across all datasets. For instance, we set the prompt length to 10 for the VPT method. To avoid adversarial overfitting [40], we apply early stopping with the epoch that has the best performance on the validation set. During adversarial finetuning, we utilize the PGD-7 attack with ${\varepsilon}=8/255$ and a step size of ${\alpha}=2/255$. When reporting test accuracy and robustness, we evaluate the models under a PGD-10 attack with ${\varepsilon}=8/255$. Additionally, we report the performance under AutoAttack (AA), which is an ensemble attack and provides a more reliable approach to evaluate the adversarial robustness. We follow the standard AutoAttack setting, i.e., untargeted APGDCE, targeted APGD-DLR, targeted FAB [9], and Square Attack [1].
@@
Methods CIFAR10 CIFAR100 Caltech256 CUB200 Dogs Avg
RanLI - Full-FT 54.9 31.8 62.6 49.7 42.5 48.3
RanLI - Adapter 50.8 29.2 66.7 40.3 68.5 51.1
RanLI - LoRa 53.0 29.0 62.8 44.7 63.3 50.6
RanLI - Bias 52.3 31.0 65.8 46.2 67.3 52.5
RanLI - VPT 52.3 29.7 66.3 40.3 67.3 51.2
RanLI - Linear 26.6 18.4 65.7 40.4 73.5 44.9
RoLI - Full-FT 56.3 32.7 70.4 52.1 73.5 57.0
RoLI - Adapter 51.1 29.1 65.8 40.6 75.2 52.4
RoLI - LoRa 53.9 30.0 65.7 44.6 74.3 53.7
RoLI - Bias 53.9 29.8 65.7 43.5 74.8 53.5
Figure 8: RoLI mitigates overfitting. When the linear head is randomly initialized (RanLI), increasing validation loss indicates that adversarial Adapter suffers from overfitting. For RoLI, the loss for the adversarial linear probing stage and Adapter finetuning stage is denoted as red and light red lines, respectively. The decreasing validation loss demonstrates that RoLI can mitigate overfitting.
@@
We extend our evaluation to Stanford Dogs using ${\varepsilon}=8/255$ and ${\varepsilon}=4/255$, as illustrated in Tab. 3. Interestingly, we observe that adversarial linear probing outperforms other methods under ${\varepsilon}=4/255$ attack but underperforms when tested against ${\varepsilon}=8/255$ adversarial attacks. We conjecture that the pretrained model lacks robustness against ${\varepsilon}=8/255$ attacks as it is trained using ${\varepsilon}=4/255$ adversarial attacks. This performance disparity confirms that adversarial linear probing inherently derives robustness from pretraining, whereas other methods predominantly achieve robustness from adversarial finetuning. Notably, RoLI consistently improves adversarial robustness even against ${\varepsilon}=8/255$ attacks, further validating the importance of a robust linear initialization for adversarial transfer learning.
@@ 6 Discussion
RoLI Mitigates Overfitting.
@@ 7 Conclusion
This paper systematically investigates how to achieve adversarial robustness in downstream tasks. We observe that PEFT methods either fail to be adversarially robust or exhibit significantly degraded adversarial robustness when initialized with a non-robust pretraining, highlighting the necessity of an adversarially robust pretraining. Given a robust pretrained model, we propose to use robust linear initialization (RoLI) followed by adversarial full finetuning or PEFT methods to achieve the best performance. This is motivated by our understanding that adversarial linear probing maximally inherits the robustness from the pretraining and the following adversarial finetuning can adapt the features to the downstream tasks effectively.
We demonstrate that RoLI outperforms random linear initialization across five image classification tasks. We hope the insights from this study greatly advance research aimed at enhancing adversarial robustness in downstream tasks. Furthermore, we believe that these findings offer invaluable guidance for developing robust models in the real world.
@@ A. Experimental Details
Table 8: Numerical results for Fig. 5.
B. Numerical Results
For Fig. 1, the numerical results are in Tab. 2 average column. The results for TWINS-AT [32] and AutoLoRa [46] are from their original papers.
We show the numerical results for Fig. 3 in Tab. 7. We do not include the CLIP results for Swin Transformer because the CLIP pretrained weights are only available for ResNet-50 [16] and ViT [14]. Tab. 8 displays the numerical results corresponding to Fig. 5.

### v2

@@ Abstract
With the prevalence of the Pretraining-Finetuning paradigm in transfer learning, the robustness of downstream tasks has become a critical concern. In this work, we delve into adversarial robustness in transfer learning and reveal the critical role of initialization, including both the pretrained model and the linear head. First, we discover the necessity of an adversarially robust pretrained model. Specifically, we reveal that with a standard pretrained model, Parameter-Efficient Finetuning (PEFT) methods either fail to be adversarially robust or continue to exhibit significantly degraded adversarial robustness on downstream tasks, even with adversarial training during finetuning. Leveraging a robust pretrained model, surprisingly, we observe that a simple linear probing can outperform full finetuning and other PEFT methods with random initialization on certain datasets. We further identify that linear probing excels in preserving robustness from the robust pretraining. Based on this, we propose Robust Linear Initialization (RoLI) for adversarial finetuning, which initializes the linear head with the weights obtained by adversarial linear probing to maximally inherit the robustness from pretraining. Across five different image classification datasets, we demonstrate the effectiveness of RoLI and achieve new state-of-the-art results. Our code is available at https://github.com/DongXzz/RoLI.
@@ 1 Introduction
In this work, we comprehensively study six popular finetuning methods in adversarial transfer learning, as shown in Fig. 2. We discover that initialization, which includes 1) the backbone initialization (i.e., a pretrained model), and 2) the initialization of the linear head that adapts features to the target domain, plays a critical role in adversarial transfer learning. We start by comparing the performance of adversarial finetuning when initialized with either a robust or a standard pretrained model. Surprisingly, all PEFT methods fail or exhibit significantly inferior performance when initialized with a standard pretrained model compared to being initialized with a robust pretrained model. Furthermore, we observe that a large-scale pretrained model, such as CLIP [37], cannot alleviate this phenomenon, highlighting the critical role of robust pretraining in adversarial transfer learning.
Given a robust pretrained model, surprisingly, we discover that adversarial linear probing outperforms other methods on certain datasets, e.g., Caltech256 [15] and Stanford Dogs [23]. We further demonstrate that this is mainly because linear probing excels in inheriting robustness from pretraining. Moreover, we identify the ability to adapt features to the specific target domain as a key factor influencing linear probing’s success, supported by a strong correlation between transferred accuracy and robustness. Based on these insights, we introduce Robust Linear Initialization (RoLI), which initializes the model’s linear head with the weights obtained from adversarial linear probing. With RoLI followed by adversarial finetuning methods, we maximize the inherited robustness along with a strong feature adaptation ability, resulting in improved performance. In summary, our contributions are as follows:
@@ 5 Initialization Matters For Finetuning
Table 1: PEFT methods, excluding Linear, demonstrate strong performance on average across CIFAR10 and Caltech256 datasets. With adversarial finetuning, PEFT methods outperform Full-FT and Linear with RanLI. Full results are available in Tab. 9
@@ 5.1 Random Linear Initialization
To answer this question, we first evaluate the adversarial robustness of six adversarial finetuning techniques. To adapt the robust pretrained model to the downstream tasks, we initialize the linear head with Random Linear Initialization (RanLI) following [42]. We perform adversarial finetuning across two model architectures, e.g., ViT and Swin, on the CIFAR10 and Caltech256 datasets. Additional details regarding the training process can be found in Sec. 5.4 and in supplementary A. Experimental Details
@@ 5.4 Robust Linear Initialization (RoLI)
Methods CIFAR10 CIFAR100 Caltech256 CUB200 Dogs Avg
RanLI - Full-FT 54.9 31.8 62.6 49.7 42.5 48.3
RanLI - Adapter 50.8 29.2 66.7 40.3 68.5 51.1
RanLI - LoRa 53.0 29.0 62.8 44.7 63.3 50.6
RanLI - Bias 52.3 31.0 65.8 46.2 67.3 52.5
RanLI - VPT 52.3 29.7 66.3 40.3 67.3 51.2
RanLI - Linear 26.6 18.4 65.7 40.4 73.5 44.9
RoLI - Full-FT 56.3 32.7 70.4 52.1 73.5 57.0
RoLI - Adapter 51.1 29.1 65.8 40.6 75.2 52.4
RoLI - LoRa 53.9 30.0 65.7 44.6 74.3 53.7
RoLI - Bias 53.9 29.8 65.7 43.5 74.8 53.5
@@
In our experiments, we use five datasets: CIFAR10 [26], CIFAR100 [26], Caltech256 [15], Caltech-UCSD Birds-200-2011 (CUB200) [45], and Stanford Dogs (Dogs) [23]. For low-resolution image datasets like CIFAR10 and CIFAR100, we resize the image to match the model input shape. For high-resolution image datasets, we apply random resizing and cropping in training and center crop in testing. It’s important to note that we integrate the resizing and normalization process directly into the model since the attacker does not have access to pre-processed images. Additional details are provided in the supplementary A. Experimental Details.
@@ Methods.
Training Details.
We use the same optimizer, AdamW [33], and a cosine scheduler with warm-up for all finetuning methods. To determine the optimal learning rate and weight decay values, we conduct a grid search, and the ranges of hyper-parameters along with the optimal combinations can be found in the supplementary A. Experimental Details. We keep method-specific hyper-parameters constant across all datasets. For instance, we set the prompt length to 10 for the VPT method. To avoid adversarial overfitting [40], we apply early stopping with the epoch that has the best performance on the validation set. During adversarial finetuning, we utilize the PGD-7 attack with ${\varepsilon}=8/255$ and a step size of ${\alpha}=2/255$. When reporting test accuracy and robustness, we evaluate the models under a PGD-10 attack with ${\varepsilon}=8/255$. Additionally, we report the performance under AutoAttack (AA), which is an ensemble attack and provides a more reliable approach to evaluate the adversarial robustness. We follow the standard AutoAttack setting, i.e., untargeted APGDCE, targeted APGD-DLR, targeted FAB [9], and Square Attack [1].
@@
We extend our evaluation to Stanford Dogs using ${\varepsilon}=8/255$ and ${\varepsilon}=4/255$, as illustrated in Tab. 4. Interestingly, we observe that adversarial linear probing outperforms other methods under ${\varepsilon}=4/255$ attack but underperforms when tested against ${\varepsilon}=8/255$ adversarial attacks. We conjecture that the pretrained model lacks robustness against ${\varepsilon}=8/255$ attacks as it is trained using ${\varepsilon}=4/255$ adversarial attacks.
This performance disparity confirms that adversarial linear probing inherently derives robustness from pretraining, whereas other methods predominantly achieve robustness from adversarial finetuning. Notably, RoLI consistently improves adversarial robustness even against ${\varepsilon}=8/255$ attacks, further validating the importance of a robust linear initialization for adversarial transfer learning.
@@ 6 Discussion
Comparison between RoLI and StdLI.
Standard Linear Initialization (StdLI) refers to the use of a simple linear probing to initialize the head, while RoLI uses adversarial linear probing. In Tab. 5, we compare the performance of RoLI and StdLI on CIFAR-10 and Caltech-256. We observe that StdLI exhibits lower performance compared to RoLI, with a robustness decrease of 2.48 across three datasets, highlighting the effectiveness of RoLI.
@@
StdLI - CIFAR-10 Caltech-256
Clean PGD Clean PGD
Full-FT 93.34(-0.84) 59.34(-1.51) 83.23(-0.55) 52.85(-2.57)
Bias 93.34(-0.04) 58.12(-0.27) 82.93(+1.15) 52.93(-1.56)
Adapter 90.08(-0.23) 55.19(-0.50) 82.72(+0.98) 53.01(-1.31)
LoRa 90.71(-1.52) 58.17(-0.45) 82.87(+1.08) 51.47(-3.03)
Table 5: StdLI is less robust than RoLI. The values in parentheses denote the performance gap (StdLI - RoLI).
@@ 7 Conclusion
This paper systematically investigates how to achieve adversarial robustness in downstream tasks. We highlight the necessity of an adversarially robust pretraining. Given a robust pretrained model, we propose to use robust linear initialization (RoLI) followed by adversarial full finetuning or PEFT methods to achieve the best performance. We demonstrate that RoLI outperforms random linear initialization across five image classification tasks. We hope the insights from this study greatly advance research aimed at enhancing adversarial robustness in downstream tasks.
Acknowledgment. We thank Chenhe Gu and Xuan Yang for the valuable discussion and insightful feedback.
@@ A. Experimental Details
Table 9: Numerical results for Fig. 4 and Tab. 1.
B. More Ablation Studies
RoLI Mitigates Overfitting.
Figure 7: RoLI mitigates overfitting. For RoLI, the loss for the adversarial linear probing stage and Adapter finetuning stage is denoted as red and light red lines, respectively. The decreasing validation loss demonstrates that RoLI can mitigate overfitting.
Figure 8: A robust pretrained model exhibits more semantically structured gradients compared to a standard pretrained model.
Image Gradient Visualization.
C. RoLI Pseudo Code
Algorithm 1 RoLI with adversarial finetuning
  Input: Dataset $D$, Model $\theta$, Parameters in linear head $\theta_{head}$, Tunable parameters in each adversarial finetuning method $\theta_{ft}$, Perturbation bound $\varepsilon$, Loss function $\mathcal{L}$
Output: Adversarially robust model $\theta^{*}$
// Initialize tunable parameters in linear head.
Random initialize $\theta_{head}$
// Robust Linear Initialization
$\hat{\theta}\leftarrow\theta_{head}$, $\theta_{frozen}\leftarrow\theta\setminus\theta_{head}$
$\hat{\theta}\leftarrow\min_{\hat{\theta}}\mathbb{E}_{(x,y)\sim D}\left[\max_{\|\delta\|_{\infty}\leq\varepsilon}\mathcal{L}(x+\delta,y;\hat{\theta}\cup\theta_{frozen})\right]$
$\theta\leftarrow\hat{\theta}\cup\theta_{frozen}$
// Adversarial FT, tunable parameters shown in Fig. 1
$\hat{\theta}\leftarrow\theta_{ft}$, $\theta_{frozen}\leftarrow\theta\setminus\theta_{ft}$
$\hat{\theta}\leftarrow\min_{\hat{\theta}}\mathbb{E}_{(x,y)\sim D}\left[\max_{\|\delta\|_{\infty}\leq\varepsilon}\mathcal{L}(x+\delta,y;\hat{\theta}\cup\theta_{frozen})\right]$
$\theta^{*}\leftarrow\hat{\theta}\cup\theta_{frozen}$
D. Numerical Results
For Fig. 1, the numerical results are in Tab. 2 average column. The results for TWINS-AT [32] and AutoLoRa [46] are from their original papers. We show the numerical results for Fig. 3 in Tab. 8. We do not include the CLIP results for Swin Transformer because the CLIP pretrained weights are only available for ResNet-50 [16] and ViT [14]. Tab. 9 displays the numerical results corresponding to Fig. 4 and Tab. 1.